# A Two-Step Concept-Based Approach for Enhanced Interpretability and Trust in Skin Lesion Diagnosis

### v1

@@ Abstract
The main challenges hindering the adoption of deep learning-based systems in clinical settings are the scarcity of annotated data and the lack of interpretability and trust in these systems. Concept Bottleneck Models (CBMs) offer inherent interpretability by constraining the final disease prediction on a set of human-understandable concepts. However, this inherent interpretability comes at the cost of greater annotation burden. Additionally, adding new concepts requires retraining the entire system. In this work, we introduce a novel two-step methodology that addresses both of these challenges. By simulating the two stages of a CBM, we utilize a pretrained Vision Language Model (VLM) to automatically predict clinical concepts, and a Large Language Model (LLM) to generate disease diagnoses based on the predicted concepts. We validate our approach on three skin lesion datasets, demonstrating that it outperforms traditional CBMs and state-of-the-art explainable methods, all without requiring any training and utilizing only a few annotated examples. The code is available at https://github.com/CristianoPatricio/2-step-concept-based-skin-diagnosis.
@@ 1 Introduction
In this work, we address the limitations of CBMs by proposing a novel two-step approach that provides concept-based explanations and generates disease diagnoses grounded in predicted concepts, all without the need for additional training. In the first stage, we utilize pretrained Vision-Language Models (VLMs) to predict the presence of a set of clinical concepts. In the second stage, these predicted concepts are incorporated into a custom designed prompt, which is used to query an LLM for a final diagnosis class to ensure the diagnosis is based on the identified concepts. Figure 1 illustrates our methodology compared to the traditional approach where a linear layer is trained to predict the final target class. Similar to CBMs, our approach allows for test-time intervention, enabling clinicians to correct concept predictions at inference time. However, unlike CBMs, our approach does not require training to provide the final diagnosis class and can be easily adapted to incorporate new concepts. We validate our approach on three well-known skin lesion datasets, demonstrating that it outperforms traditional CBMs and state-of-the-art explainable methods, all without requiring any training and using only a few annotated examples within the prompt. Our contributions are as follows:
  • We introduce a novel two-step approach for concept-based explainability in skin lesion diagnosis that can provide the final diagnosis class without the need for additional training;
  • We reduce the annotation burden associated with CBMs by leveraging the zero-shot capabilities of pretrained Vision-Language Models (VLMs) to automatically predict the presence of concepts;
  • We provide the final diagnosis class through the use of a tailored prompt and using the few-shot capabilities of LLMs;
  • Our approach demonstrates superior interpretability and performance compared to traditional CBMs across three public benckmarks.
Figure 1: Overview of the proposed framework. The linear classifier layer (left) is replaced by a Large Language Model (LLM) (right), which grounds its responses on clinical concepts predicted by a pretrained vision-language model (VLM). This approach is training-free and not restricted by predefined labels, allowing the LLM to generate diverse diagnostic possibilities for different diseases.
@@ 4 Experiments
In this section, we begin by describing our experimental setup, including details of the models, datasets, and evaluation metrics used. Following that, we present the main results of our approach, assessing both zero-shot and few-shot performance based on Balanced Accuracy (BACC), Sensitivity, and Specificity.
@@ 4.2 Implementation Details
For the experiments with concept-based models, we reproduce the results using the official implementation of CBM [4]. For the pretrained VLMs we use the checkpoints from the HuggingFace library for CLIP, BiomedCLIP and MONET. For CBI-VLM and ExpLICD we use their official implementation, available in [17] and [19], respectively.
@@
For the selection strategy used to perform ICL, the image similarity is determined using the cosine similarity between image features, which are extracted from the visual encoder of ExpLICD. Further details and documentation are available at https://github.com/CristianoPatricio/2-step-concept-based-skin-diagnosis. All experiments are run on a single A40 GPU with 48GB of VRAM.
5 Results
We compare the performance of our approach (VLM + LLM) against two baseline approaches:
@@ 5 Results
Contrary to these methods, our approach (VLM + LLM) does not require training at all. We predict concepts using a pretrained VLM and prompt an LLM to generate the final diagnosis.
@@ 5.1 Disease Classification Performance
Model Trainable Dataset PH2 Derm7pt HAM10000 Sens. Spec. BAcc Sens. Spec. BAcc Sens. Spec. BAcc Image to Concepts to Label Classification $[x\rightarrow c\rightarrow y]$ CBM [4] 42.04 98.06 70.14 69.31 87.67 79.03 51.00 94.00 70.83 VLM+ Linear Classifier CLIP [13] + Linear Classifier 59.29 96.32 77.80 37.62 91.32 64.47 21.00 94.00 57.50 BiomedCLIP [14] + Linear Classifier 22.78 98.18 60.49 6.93 99.09 53.01 0.00 100.0 50.00 MONET [18] + Linear Classifier 71.43 96.79 84.11 53.47 92.69 73.08 20.00 92.00 56.00 CBI-VLM [17] + Linear Classifier 24.60 98.22 61.41 50.50 89.95 70.22 56.00 92.00 74.00 ExpLICD [19] + Linear Classifier 58.49 96.47 77.48 55.45 95.43 75.44 60.00 87.00 73.50 Ours (VLM + LLM) CLIP [13]    + MMed-Llama3-8B-EnIns 88.17 3.71 45.94 59.41 14.61 37.01 87.00 2.00 44.50    + Mistral-7B-Instruct 8.33 96.90 52.62 19.80 88.58 54.19 11.00 94.00 55.00 BiomedCLIP [14]    + MMed-Llama3-8B-EnIns 21.75 87.38 54.56 21.78 70.78 46.28 21.00 70.00 45.50    + Mistral-7B-Instruct 5.00 93.59 49.29 9.90 79.00 44.45 10.00 78.00 44.00 MONET [18]    + MMed-Llama3-8B-EnIns 100.0 28.67 64.33 89.11 31.96 60.54 77.00 45.00 61.00    + Mistral-7B-Instruct 73.17 61.30 67.24 73.27 62.56 67.91 40.00 67.00 53.50 CBI-VLM [17]    + MMed-Llama3-8B-EnIns 6.11 96.27 51.19 6.93 97.26 52.10 77.00 11.00 44.00    + Mistral-7B-Instruct 86.67 0.59 43.63 99.01 11.87 55.44 100.0 0.00 50.00 ExpLICD [19]    + MMed-Llama3-8B-EnIns 58.49 97.65 78.07 65.35 91.78 78.56 66.00 86.00 76.00    + Mistral-7B-Instruct 72.06 82.82 77.44 67.33 90.87 79.10 72.00 79.00 75.50    + GPT-4 72.78 93.96 83.37 67.33 90.87 79.10 77.00 72.00 74.50 🟊$\bigstar$ (ExpLICD + LLM w/ few-shot) 79.44 90.65 85.05 67.33 92.24 79.78 76.00 74.00 75.00
@@
Our proposed approach consistently outperforms all baseline methods across all datasets. In contrast, CLIP-like models, such as CLIP and BiomedCLIP, generally show poor zero-shot performance, especially in terms of sensitivity, as they frequently predict the benign class (nevus). On the other hand, Vision-Language Models (VLMs) fine-tuned on dermoscopic images perform significantly better, with MONET and ExpLICD showing the highest accuracy. Although CBI-VLM was also fine-tuned on dermoscopic data, it underperformed compared to MONET and ExpLICD, likely due to its reliance on image-text alignment only in the final layers of CLIP, trained on generic data, while keeping the remaining weights frozen. However, when combining ExpLICD with MMed for PH2 and HAM10000, and ExpLICD with Mistral for Derm7pt, our method surpasses all baselines. This can be attributed to ExpLICD’s use of a pretrained vision-language model and its incorporation of textual diagnostic criteria as knowledge anchors. These anchors are aligned with predefined visual concept tokens through contrastive loss, resulting in more fine-grained alignment than traditional image-text alignment in VLMs. We also evaluate the combination of ExpLICD + GPT-4, motivated by the strong performance of this VLM with MMed and Mistral. However, no substantial improvements were observed compared to using open-source LLMs like MMed and Mistral.
Furthermore, we show that the results for PH2 and Derm7pt can be further improved by incorporating additional context into the prompt through the inclusion of demonstration examples, selected using our retrieval strategy, as shown in Figure 3.
@@
As shown in Figure 3, for PH2, we increase BAcc by 6% and 7% when using Mistral and MMed, respectively, with only one demonstration example. In the case of Derm7pt, the zero-shot performance improves by 0.68% when using $K=4$ demonstrations in Mistral and 0.42% with $K=8$ demonstration examples in MMed. However, for HAM10000, no performance improvement was observed when using In-Context Learning (ICL).
@@ 5.2 Ablation study on disease classification
To further validate the effectiveness of our method, we conduct ablation studies on the disease classification strategy. Specifically, we explore direct label prediction from the input image ($x\rightarrow y$), discarding the intermediate concept prediction step. As shown in Table 4, directly comparing image features with encoded class labels yields lower performance than our proposed method (last row of Table 4), which first predict the clinical concepts and then uses that information for class label prediction. Nonetheless, CBI-VLM achieves the highest performance across most datasets, benefiting from fine-tuning that aligns image and label features within the same class. These results emphasize the value of an intermediate concept prediction step, enhancing both performance and interpretability.
@@ 7 Conclusions and Future Work
We present a novel two-step methodology that utilizes off-the-shelf Vision-Language Models (VLMs) to predict clinical concepts and a Large Language Model (LLM) to generate disease diagnoses based on those concepts. This approach enhances transparency and explainability in decision-making, which is crucial for high-stakes applications like medical use cases. Our evaluation on three skin lesion datasets shows that our method outperforms traditional Concept Bottleneck Models (CBMs) and state-of-the-art explainable approaches, without requiring additional training and using only a few demonstration examples, thereby addressing the two main limitations of CBMs. Furthermore, our approach easily incorporates new concepts without retraining and supports test-time interventions. While our methodology was evaluated only on skin image datasets, it is generalizable and can be applied to other domains. In future work, we plan to explore the integration of visual explanations to further enhance trust and promote the adoption of automated diagnosis systems in clinical settings.
Acknowledgments    This work was funded by the Portuguese Foundation for Science and Technology (FCT) under the PhD grant “2022.11566.BD”, and supported by NOVA LINCS (UIDB/04516/2020) with the financial support of FCT.IP.

### v2

@@ Abstract
The main challenges hindering the adoption of deep learning-based systems in clinical settings are the scarcity of annotated data and the lack of interpretability and trust in these systems. Concept Bottleneck Models (CBMs) offer inherent interpretability by constraining the final disease prediction on a set of human-understandable concepts. However, this inherent interpretability comes at the cost of greater annotation burden. Additionally, adding new concepts requires retraining the entire system. In this work, we introduce a novel two-step methodology that addresses both of these challenges. By simulating the two stages of a CBM, we utilize a pretrained Vision Language Model (VLM) to automatically predict clinical concepts, and an off-the-shelf Large Language Model (LLM) to generate disease diagnoses based on the predicted concepts. Furthermore, our approach supports test-time human intervention, enabling corrections to predicted concepts, which improves final diagnoses and enhances transparency in decision-making. We validate our approach on three skin lesion datasets, demonstrating that it outperforms traditional CBMs and state-of-the-art explainable methods, all without requiring any training and utilizing only a few annotated examples. The code is available at https://github.com/CristianoPatricio/2-step-concept-based-skin-diagnosis.
@@ 1 Introduction
In this work, we address the limitations of CBMs by proposing a novel two-step approach that provides concept-based explanations and generates disease diagnoses grounded in predicted concepts, all without the need for additional training. In the first stage, we utilize pretrained Vision-Language Models (VLMs) to predict the presence of a set of clinical concepts. In the second stage, these predicted concepts are incorporated into a custom designed prompt, which is used to query an LLM for a final diagnosis class to ensure the diagnosis is based on the identified concepts. Figure 1 illustrates our methodology compared to the traditional approach where a linear layer is trained to predict the final target class. Similar to CBMs, our approach allows for test-time intervention, enabling clinicians to correct concept predictions at inference time. However, since the predicted concepts by VLM are presented as plain text for the final diagnosis, this enables more intuitive human intervention compared to adjusting numerical values in the bottleneck layer, as in traditional CBM test-time interventions. Additionally, leveraging a VLM to predict concepts offers the advantage of incorporating both visual and textual information during model inference. In the other hand, unlike CBMs, our approach does not require training to provide the final diagnosis class and can be easily adapted to incorporate new concepts. We validate our approach on three well-known skin lesion datasets, demonstrating that it outperforms traditional CBMs and state-of-the-art explainable methods, all without requiring any training and using only a few annotated examples within the prompt. Our contributions are as follows:
  • We introduce a novel two-step approach for concept-based explainability in skin lesion diagnosis without requiring additional training;
  • We reduce the annotation burden of CBMs by leveraging zero-shot capabilities of pretrained VLMs to automatically predict clinical concepts;
  • We provide the final diagnosis class using tailored prompts and leveraging the few-shot capabilities of LLMs;
  • Our approach enables human test-time intervention, allowing corrections to any miscpredicted concepts;
  • We demonstrate superior interpretability and performance compared to traditional CBMs on three public benckmarks.
The remainder of the paper is organized as follows: Section 2 reviews related work in the context of our study. Section 3 introduces the proposed two-step methodology. Section 4 outlines the experimental setup and implementation details. Section 5 presents key results and analysis. A general discussion is provided in Section 6. Finally, Section 7 and Section 8 discuss the limitations and conclude the paper, respectively.
Figure 1: Overview of the proposed framework. The linear classifier layer (left) is replaced by a pretrained Large Language Model (LLM) (right), which grounds its responses on clinical concepts predicted by a pretrained vision-language model (VLM). This approach is training-free and not restricted by predefined labels, allowing the LLM to generate diverse diagnostic possibilities for different diseases.
@@ 4 Experiments
In this section, we outline our experimental setup, including details of the adopted models, datasets, and evaluation metrics, as well as the implementation details.
@@ 4.2 Implementation Details
For experiments with concept-based models, we reproduce the results using the official implementation of CBM [4]. For pretrained VLMs, we use the checkpoints from the HuggingFace library for CLIP, BiomedCLIP and MONET. Additionally, we employ the official implementations of CBI-VLM and ExpLICD, as provided in [17] and [19], respectively. Checkpoints for Mistral-7B-Instruct and MMed-LLama3-8B-EnIns are also sourced from the HuggingFace library.
@@
For the selection strategy used to perform ICL, the image similarity is determined using the cosine similarity between image features, which are extracted from the visual encoder of ExpLICD. Further details, documentation and a demo of our approach are available at https://github.com/CristianoPatricio/2-step-concept-based-skin-diagnosis. All experiments are run on a single A40 GPU with 48GB of VRAM.
5 Results and Analysis
This section presents the key results of our two-step approach for skin lesion classification across three dermoscopic datasets. Specifically, we compare the performance of our approach (VLM + LLM) against two baseline approaches:
@@ 5 Results and Analysis
Our two-step approach, VLM + LLM, leverages a pretrained VLM to predict concepts from an input image (Step 1) and prompts an off-the-shelf LLM to generate the final diagnosis (Step 2) based on the predicted concepts.
@@ 5.1 Disease Classification Performance
Model Trainable Dataset PH2 Derm7pt HAM10000 Sens. Spec. BAcc Sens. Spec. BAcc Sens. Spec. BAcc Image to Concepts to Label Classification $[x\rightarrow c\rightarrow y]$ CBM [4] 42.04 98.06 70.14 69.31 89.50 77.56 51.00 94.00 70.83 VLM+ Linear Classifier CLIP [13] + Linear Classifier 59.29 96.32 77.80 37.62 91.32 64.47 21.00 94.00 57.50 BiomedCLIP [14] + Linear Classifier 22.78 98.18 60.49 6.93 99.09 53.01 0.00 100.0 50.00 MONET [18] + Linear Classifier 71.43 96.79 84.11 53.47 92.69 73.08 20.00 92.00 56.00 CBI-VLM [17] + Linear Classifier 24.60 98.22 61.41 50.50 89.95 70.22 56.00 92.00 74.00 ExpLICD [19] + Linear Classifier 58.49 96.47 77.48 55.45 95.43 75.44 60.00 87.00 73.50 Ours (VLM + LLM) CLIP [13]    + MMed-Llama3-8B-EnIns 88.17 3.71 45.94 59.41 14.61 37.01 87.00 2.00 44.50    + Mistral-7B-Instruct 8.33 96.90 52.62 19.80 88.58 54.19 11.00 94.00 55.00 BiomedCLIP [14]    + MMed-Llama3-8B-EnIns 21.75 87.38 54.56 21.78 70.78 46.28 21.00 70.00 45.50    + Mistral-7B-Instruct 5.00 93.59 49.29 9.90 79.00 44.45 10.00 78.00 44.00 MONET [18]    + MMed-Llama3-8B-EnIns 100.0 28.67 64.33 89.11 31.96 60.54 77.00 45.00 61.00    + Mistral-7B-Instruct 73.17 61.30 67.24 73.27 62.56 67.91 40.00 67.00 53.50 CBI-VLM [17]    + MMed-Llama3-8B-EnIns 6.11 96.27 51.19 6.93 97.26 52.10 77.00 11.00 44.00    + Mistral-7B-Instruct 86.67 0.59 43.63 99.01 11.87 55.44 100.0 0.00 50.00 ExpLICD [19]    + MMed-Llama3-8B-EnIns 58.49 97.65 78.07 65.35 91.78 78.56 66.00 86.00 76.00    + Mistral-7B-Instruct 72.06 82.82 77.44 67.33 90.87 79.10 72.00 79.00 75.50    + GPT-4 72.78 93.96 83.37 67.33 90.87 79.10 77.00 72.00 74.50 🟊$\bigstar$ (ExpLICD + LLM w/ few-shot) 79.44 90.65 85.05 67.33 92.24 79.78 76.00 74.00 75.00
@@
Zero-shot classification performance
Using our two-step methodology (VLM + LLM), CLIP-like models such as CLIP and BiomedCLIP generally exhibit poor zero-shot performance, especially in terms of sensitivity, as they frequently predict the benign class (nevus). In contrast, Vision-Language Models (VLMs) fine-tuned on dermoscopic images perform significantly better, with MONET and ExpLICD showing the highest accuracy. Although CBI-VLM was also fine-tuned on dermoscopic data, it underperformed compared to MONET and ExpLICD. This may be due to its reliance on image-text alignment only in the final layers of CLIP, which were trained on generic data, while keeping the remaining weights frozen. When combining ExpLICD with MMed for PH2 and HAM10000, and ExpLICD with Mistral for Derm7pt, we increase BAcc over traditional CBM: 8% on PH2, 1.5% on Derm7pt, and 5% on HAM10000. These gains can be attributed to ExpLICD’s use of a pretrained vision-language model and its incorporation of textual diagnostic criteria as knowledge anchors. These anchors are aligned with predefined visual concept tokens through contrastive loss, resulting in more fine-grained alignment than traditional image-text alignment in VLMs. We also evaluate the combination of ExpLICD + GPT-4 due the strong performance of ExpLICD with MMed and Mistral. However, no substantial improvements were observed compared to using open-source LLMs like MMed and Mistral.
Few-shot classification performance
Figure 3 shows that the results for PH2 and Derm7pt can be further improved by incorporating additional context into the prompt through the inclusion of $n$ demonstration examples, selected using our retrieval strategy. As shown in Figure 3, for PH2, we achieve a score of 83.56% and 85.05% when using Mistral and MMed, respectively, with only one demonstration example, significantly outperforming the zero-shot performance. For Derm7pt, the zero-shot performance improves with $K=4$ demonstrations in Mistral and with $K=8$ demonstration examples in MMed. Although no BAcc improvement was observed for HAM10000 using ICL, we achieved increased sensitivity with two demonstration examples in MMed and one in Mistral.
@@ 5.2 Ablation study on disease classification
To further validate the effectiveness of our method, we conduct ablation studies on the disease classification strategy. Specifically, we explore direct label prediction from the input image ($x\rightarrow y$) by discarding the intermediate concept prediction step. Instead, we compare image features to short text prompt describing the class, detailed in Table 4.
Model Template
BiomedCLIP This is a dermoscopic image of {class}
CLIP This is a dermoscopic image of {class}
MONET This is skin image of {class}
CBI-VLM This is dermatoscopy of {class}
ExpLICD This is a dermoscopic image of {class}
Table 4: Prompts used for zero-shot image classification. The {class} is replaced by “Nevus” or “Melanoma” in our experiments.
As shown in Table 5, directly comparing image features with encoded class labels yields lower performance than our proposed method (last row of Table 5), which first predict the clinical concepts and then uses that information for class label prediction. Nonetheless, CBI-VLM achieves the highest performance across most datasets, benefiting from fine-tuning that aligns image and label features within the same class. These results emphasize the value of an intermediate concept prediction step, enhancing both performance and interpretability.
@@ 5.4 Diagnostic Interpretability
6 General Discussion
The effect of few-shot prompting in the overall performance
Our proposed two-step approach consistently outperforms all baseline methods across datasets when using ICL, demonstrating the utility of incorporating additional context into prompts with a few demonstration examples to facilitate LLM reasoning. This improvement was particularly noticeable in the PH2 and Derm7pt datasets. However, no performance gains were observed for HAM10000, which can be attributed to two key factors:
  • Incorrect prediction of concepts: Our empirical analysis shows that predicted concepts have a significant impact on LLM responses. While demonstration examples can guide the LLM to provide correct diagnoses, situations may arise where incorrectly predicted concepts, such as predicting features commonly associated with melanomas when the ground truth label is nevus, lead the LLM to produce incorrect disease diagnoses, as its response is grounded in the predicted concepts.
  • Inaccuracy in demonstration examples: Another observed issue is when demonstration examples contain inaccurate associations between concepts and ground truth labels. Since the concepts are automatically predicted by the VLM, there is no guarantee that demonstration examples from the training set accurately represent the correct associations. This mismatch could mislead the LLM into providing incorrect diagnoses.
Overall, these issues were frequently observed in HAM10000 dataset, suggesting that it contains more challenging images, which may contribute to a higher False Positive rate. Further refinement of concept prediction and selection of high-quality demonstration examples could mitigate these limitations.
Transparency and interpretability in model responses
Overall, the results presented in Section 5 highlight the effectiveness of our two-step approach in enabling pretrained VLMs to predict clinical concepts from images and use those concepts to ground the final diagnosis. This enhances both interpretability and transparency in the model’s responses, all without requiring additional training and using only a few annotated examples. Notably, our approach outperforms traditional CBMs and state-of-the-art XAI methods across all datasets.
Human interaction through test-time intervention
Our two-step methodology allows for human intervention in the clinical concepts predicted by the VLM. Since the concepts are presented in plain text to the user, this offers a significant advantage over traditional CBMs, which require adjusting numerical values in the bottleneck layer. Furthermore, our results in Section LABEL:subsec:tti underscore the importance of this test-time intervention in enhancing overall performance.
Overall advantages of the proposed two-step approach
When comparing our work with studies [9, 10, 11, 12] using similar techniques, our two-step approach stands out for multiple reasons:
  • Training-free methodology: We use pretrained models with medical knowledge that require only two inference steps (taking less than 3 seconds) to generate an interpretable final diagnosis. In contrast, existing CBMs and supervised methods require extensive training and are annotation-dependent;
  • Intuitive test-time intervention: We enable intuitive human intervention to correct any predicted concept by editing plain text;
  • Flexible incorporation of concepts and class labels: Additional concepts or class labels can be included without the need to retrain the entire architecture. In contrast, CBMs and supervised task-specific models are limited to a predefined set of class labels.
@@ 8 Conclusions and Future Work
We present a novel two-step methodology that utilizes off-the-shelf Vision-Language Models (VLMs) to predict clinical concepts and a Large Language Model (LLM) to generate disease diagnoses based on those concepts. This approach enhances transparency and explainability in decision-making, which is crucial for high-stakes applications like medical use cases. Our evaluation on three skin lesion datasets demonstrates that our method outperforms traditional Concept Bottleneck Models (CBMs), achieving a relative improvement of 21% on PH2, 3% on Derm7pt, and 6% on HAM10000. Additionally, it surpasses state-of-the-art explainable approaches without requiring additional training and uses only a few demonstration examples, thereby addressing the two primary limitations of CBMs. Furthermore, our approach easily incorporates new concepts without retraining and supports test-time interventions. While our methodology was evaluated only on skin image datasets, it is generalizable and can be applied to other domains. In future work, we plan to explore the integration of visual explanations to further enhance trust and promote the adoption of automated diagnosis systems in clinical settings.
Acknowledgments    This work was funded by the Portuguese Foundation for Science and Technology (FCT) under the PhD grant “2022.11566.BD”, and supported by UID/04516/NOVA Laboratory for Computer Science and Informatics (NOVA LINCS) with the financial support of FCT.IP.